No Observing tonight.

A hail storm in the afternoon produce half an inch of hail accumulated on the dome. We had serious water leaks that affected the LBC Tree House. Most of the PCs got wet and both the Red Tech and Blue Sci Computers failed to restart.

After many hours of ventilating the Tree House with a fan the LBC team succeeded in resuscitating all the 4 PCs at around 10 UT. Night was lost. We could have observed for some hours though with thick cirrus clouds.

-- MarcoPedani - 12 Sep 2009
Topic revision: r1 - 12 Sep 2009, MarcoPedani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback